OpenStudy (anonymous):

Marie, Brad, and Doug went to the mall. Brad spent three times as much money as Marie. Doug spent fifteen dollars more than Brad. If Doug spent $60 at the mall, which of the following equations should be used to find out how much money Marie spent? A)3(m + 15) = 60 (where m represents the amount of money Marie spent) B)15m + 3 = 60 (where m represents the amount of money Marie spent) C)3m + 15 = 60 (where m represents the amount of money Marie spent) D)3m – 15 = 60 (where m represents the amount of money Marie spent)

OpenStudy (anonymous):

D : B + 15 = 60 (solve for B) B = 60 - 15 B = 45 B : 3M = 45 (solve for M) M = 45/3 M = 15 ANSWER : 3M + 15 = 60

OpenStudy (anonymous):

thank you @kelliegirl22

OpenStudy (anonymous):

your welcome :)